OpenStudy (anonymous):

solve the following equation 7/9*x=-21

sam (.sam.):

(7)/(9)*x=-21 7x/9=-21 7x=-189 x=-27

OpenStudy (anonymous):

\[\left[\begin{matrix}? & ? \\ ? & ?\end{matrix}\right]\]